A constant horizontal force of 5.0 N is applied to a 1.2 kg cart. The cart is initially at rest. How fast will it be moving after a time of 3.0 seconds


1
Expert's answer
2021-03-23T11:15:35-0400
"F\Delta t=m\Delta v,""F\Delta t=m(v_f-v_i)=mv_f,""v_f=\dfrac{F\Delta t}{m}=\dfrac{5.0\ N\cdot3.0\ s}{1.2\ kg}=12.5\ \dfrac{m}{s}."